Bug 164166 - BroadCom NIC on HP DL 380 and 580 maintains link but driver stops TX/RX - (tg3 and bcm5700)
BroadCom NIC on HP DL 380 and 580 maintains link but driver stops TX/RX - (tg...
Status: CLOSED NOTABUG
Product: Red Hat Enterprise Linux 3
Classification: Red Hat
Component: kernel (Show other bugs)
3.0
i386 Linux
medium Severity medium
: ---
: ---
Assigned To: John W. Linville
Brian Brock
:
Depends On:
Blocks:
  Show dependency treegraph
 
Reported: 2005-07-25 11:39 EDT by Ian McLeod
Modified: 2007-11-30 17:07 EST (History)
4 users (show)

See Also:
Fixed In Version:
Doc Type: Bug Fix
Doc Text:
Story Points: ---
Clone Of:
Environment:
Last Closed: 2005-09-15 17:04:39 EDT
Type: ---
Regression: ---
Mount Type: ---
Documentation: ---
CRM:
Verified Versions:
Category: ---
oVirt Team: ---
RHEL 7.3 requirements from Atomic Host:
Cloudforms Team: ---


Attachments (Terms of Use)

  None (edit)
Description Ian McLeod 2005-07-25 11:39:45 EDT
From Bugzilla Helper:
User-Agent: Mozilla/5.0 (X11; U; Linux i686; en-US; rv:1.7.5) Gecko/20041215 Firefox/1.0 Red Hat/1.0-12.EL4

Description of problem:
All network connectivity is lost on HP DL380 and HP DL580 servers using the built in BroadCom NIC.  Both the kernel and the switch continue to show an active link, and the RX packet counter in ifconfig continues to increment.  However, no packets are sent or received by the tg3 or bcm5700 drivers.  The problem persists until a reboot.

Problem occurs more or less randomly and cannot be reliably reproduced.  However, in a large (300+) deployment of DL 380 and DL 580s at a large client this problem manifests itself a few times a week.

Version-Release number of selected component (if applicable):


How reproducible:
Couldn't Reproduce

Steps to Reproduce:
This has been reproduced only insofar as it has reliably occured on 1-2 machines a week in a large deployment.

Additional info:
Comment 1 Ian McLeod 2005-07-25 11:44:18 EDT
HP has released a BroadCom firmware update that purports to address this
particular issue.  It is available at:

http://h20000.www2.hp.com/bizsupport/TechSupport/Document.jsp?objectID=PSD_EU050629_CW01

The syptoms described in this HP support document match the syptoms we've
observed in our environment.  We will be applying this update to a collection of
affected servers and will update this bugzilla with results of this testing.
Comment 2 Ernie Petrides 2005-07-25 14:47:01 EDT
Just read comment #1 -- let us know if the f/w update resolves this.  Thanks.
Comment 3 John W. Linville 2005-09-15 17:04:39 EDT
Closed due to lack of response.  Please re-open if/when requesting information 
becomes available...thanks! 
Comment 4 Ernie Petrides 2005-09-15 17:12:26 EDT
Since comment #1 suggested this might have been a f/w issue,
I'm changing the disposition to NOTABUG.

Note You need to log in before you can comment on or make changes to this bug.